OpenStudy (brightemmerald):

What is the vertex of the quadratic function below? y= 3x^3 - 12x +17 A. (-2, -5) B. (3, 17) C. (-3, 17) D. (2, 5)

OpenStudy (ajspeller):

I suspect that you mis-typed the equation ax^2+bx+c=0 Your a value is 3 Your b value is -12 Your c value is 17 x = -b/2a y= f(-b/2a)

OpenStudy (brightemmerald):

Here it is, I hope this makes better since.

OpenStudy (ajspeller):

Here is a tutorial to help explain the process http://www.youtube.com/watch?v=BODGlEl-zWM&t=4m6s Hope this helps!

OpenStudy (ajspeller):

sense not since

OpenStudy (danjs):

You can turn that standard form y=ax^2 + bx + c , into a vertex form y=a(x - h)^2 + k where the vertex is the point (h,k)

OpenStudy (danjs):

y = 3x^2 - 12x + 17 factor 3 from the first 2 terms y = 3*(x^2 - 4x) + 17 complete the square adding 4 and taking away 3*4=12 y = 3*(x^2 - 4x + 4) - 12 + 17 factor y = 3*(x - 2)^2 + 5 There is the vertex form, you can see the vertex is (h,k)=(2 , 5)